OpenStudy (anonymous):

Using the formula C=5/9(F-32); f=fahrenheit, if the celcius (C) temperature increased 35 degrees, by how many degrees would the Fahrenheit temperature be increased? A) 19 & 4/9 B) 31 degrees C) 51 Degrees D) 63 degrees E) 82 degrees

OpenStudy (anonymous):

\[C=\frac{5}{9} *(F-32)\] \[F=\frac{9C+160}{5}\] celcius (C) temperature increased 35 degrees \[C+35=\frac{5}{9} *(F_ {new} -32)\] \[9C+315=5(F_ {new} -32)\] \[9C+315=5F_ {new}-160\] \[F_ {new}=\frac{9C+475}{5}\] increase in Fahrenheit=\[F_ {new}-F=\frac{9C+475}{5}-\frac{9C+160}{5}\] \[F_ {new}-F=\frac{315}{5}=63\]

OpenStudy (anonymous):

Thank you